How many liters of an 11% salt solution must be mixed with 16 liters of 44% salt solution to obtain a 23% salt solution?
Answer by stanbon(57203)

How many liters of an 11% salt solution must be mixed with 16 liters of 44% salt solution to obtain a 23% salt solution?
-----------------------------
salt + salt = salt
0.11x + 0.44*16 = 0.23(x+16)
----
Multiply thru by 100 to get:
11x + 44*16 = 23x + 23*16
12x = 21*16
x = 28 liters (amt. of 11% solution needed)
==============================================
Cheers,
Stan H.
==========